# Predicting the Geolocation of Tweets Using transformer models on Customized Data

### v4

@@ 2 Related works
Generally, there are three categories of home location granularity: administrative regions, geographical grids, and geographical coordinates. As for the tweet location, point-of-interests (POIs) or coordinates are broadly adopted as representations of tweet locations, instead of administrative regions or grids.
@@
Another source of location-indicative information is based on the user interconnection graphs, also referenced as Twitter network in Twitter-related studies. [44]. When tasked with the prediction of users’ home locations through an analysis of the Twitter network, one might initially conjecture that one’s home location is very likely to coincide with their friends’ home locations. Likewise, when users mention each other or have a conversation with each other, this interaction alludes to a closer relationship or shared interests between the two users, also referenced as social closeness.
@@
Despite short in-length textual content, a tweet is accompanied by rich context like timestamps and geo-tags associated with it, as well as various attributes of its author profile. For instance, when sending out a tweet, it is automatically accompanied by a posting timestamp. Moreover, in the era of widespread GPS-enabled devices such as smartphones and tablets, users have the option to append geo-tags to their tweets, indicating their current locations. In the case of Twitter, the spatial context is represented as machine-generated geographic object descriptions with countries, cities, place types, and full names attributed by Twitter to the user-defined geolocation coordinates.
@@
Continuing the exploration of probabilistic models used in the relevant studies, Iso in the work [12] proposed the convolutional mixture density network which is applied to the tweets’ content to estimate the parameters of the GMM and employ the mode value of estimated distribution density as the predicted coordinates for tweets. The authors also claim that the choice of a loss function affects the model performance and describe a basic logic of the probabilistic model loss computation using negative log-likelihood as shown in Eq. (NLLH). Similarly, Li in the work [17] utilizes the negative log-likelihood of the GMM over all training examples. The attention score is used as the mixture weight of the Gaussian components, which is then combined to output a Gaussian mixture distribution over the geographic space. In contrast, we propose the prediction of all GMM parameters (weights, covariance) by the model itself rather than the application of available attention weights.
@@
In the study conducted by Simanjuntak [33], the task of predicting Twitter users’ home location was explored using Long-Short Term Memory (LSTM) and BERT models. The dataset used in the experiment comprised text content and user metadata context of tweets in the Indonesian language. Given that the base model has a limited input capacity of 512 tokens (words), it is not feasible to process large text corpora that comprise all user tweets.
Such constraints posed a challenge in accurately classifying a user’s home location, as concatenating all of the user’s tweets into a single text would result in an input exceeding the limit. As a result, users’ location prediction was performed on a per-tweet basis, and it was observed that the majority vote approach frequently (42%) led to misclassification of the user’s home location. Thus only the task of tweet geolocation prediction could be efficiently solved using the BERT-base models. In contrast, this work focuses on processing smaller (less than 300 words) text samples and aggregating a set of probabilistic predictions in a form of GMMs to estimate the set of most probable user’s home location points as described in Section A.
@@ 3.3.1 Worldwide evaluation metrics by model type
The Probabilistic Single/Multiple Outcome Prediction (PSOP/PMOP) models have slightly outperformed the straightforward approach of Geospatial Single/Multiple Outcome Prediction (GSOP/GMOP) models on the median error distance metrics as shown in Table 4. This observation supports findings in the work [17] that demonstrated the superiority of predicting a probability distribution instead of a single point in space, since it is possible to have several candidate locations that pull the single output in different directions. Similarly, the MOP-type architecture outperformed SOP-type in both Probabilistic and Geospatial variations, though the difference was relatively low.
@@ 4 Discussion and Conclusion
The user-level geolocation is certainly effective for some applications, however tweet-level geolocation supports fine-grained analyses and thus was chosen as the main goal of this study. Nonetheless, the problem of the user’s home location prediction was explored as one of the possible estimates based on the PMOP-type model predictions. Although this wasn’t the foremost performance measure, the results reveal a higher spatial accuracy on the per-user task than on the per-tweet measurements. Hence PMOP-type models are suggested for the user’s home location prediction as the comparable transformers-based solution.

### v5

@@ 2 Related works
According to [44], the granularity of user’s home location prediction task has three levels: administrative regions, geographical grids, and geographical coordinates. As for the separate tweet locations, coordinates or point-of-interests (POIs) are used instead of administrative regions and grids.
@@
Another source of location-indicative information is based on the user interconnection graphs, also referenced as Twitter network in Twitter-related studies. [44]. When tasked with the prediction of users’ home locations through an analysis of the Twitter network, we can assume that the user’s home is close to their friends’ home locations. Likewise, mentions and conversations allude to a closer relationship or shared interests between the two users, also referenced as social closeness.
@@
Despite short in-length textual content, a tweet is accompanied by rich context like timestamps and geo-tags associated with it, as well as various attributes of its author profile. For instance, when sending out a tweet, it is automatically accompanied by a posting timestamp. Moreover, in the era of widespread GPS-enabled devices such as smartphones and tablets, users have the option to append geo-tags to their tweets, indicating their current locations [44]. In the case of Twitter, the spatial context is represented as machine-generated geographic object descriptions with countries, cities, place types, and full names attributed by Twitter to the user-defined geolocation coordinates.
@@
Continuing the exploration of probabilistic models used in the relevant studies, Iso in the work [12] proposed the convolutional mixture density network which is applied to the tweets’ content to estimate the parameters of the GMM and employ the mode value of estimated distribution density as the predicted coordinates for tweets. The authors also claim that the choice of a loss function affects the model performance and describe a basic logic of the probabilistic model loss computation using negative log-likelihood as shown in Eq. (NLLH). Similarly, Li in the work [17] utilizes the negative log-likelihood of the GMM over all training examples. Notably, the weights of the Gaussian components were obtained directly from the model’s attention scores. In contrast, we propose the prediction of all GMM parameters (weights, covariance) by the model itself rather than the application of available attention weights.
@@
In the study conducted by Simanjuntak [33], the task of predicting Twitter users’ home location was explored using Long-Short Term Memory (LSTM) and BERT models. The authors used a dataset in the Indonesian language containing text content and user metadata context of tweets. Given that the base model has a limited input capacity of 512 tokens (words), it is not feasible to process large text corpora that comprise all user tweets.
Such constraints were challenging for classifying a user’s home location, as concatenating all of the user’s tweets into a single text would result in an input exceeding the limit. Authors state that the majority vote approach for per-tweet predictions in 42% cases resulted in misclassification of user home location. Thus only the task of tweet geolocation prediction could be efficiently solved using the BERT-base models. In contrast, this work focuses on processing smaller (less than 300 words) text samples and aggregating a set of probabilistic predictions in a form of GMMs to estimate the set of most probable user’s home location points as described in Section A.
@@ 3.3.1 Worldwide evaluation metrics by model type
The Probabilistic Single/Multiple Outcome Prediction (PSOP/PMOP) models have slightly outperformed the straightforward approach of Geospatial Single/Multiple Outcome Prediction (GSOP/GMOP) models on the median error distance metrics as shown in Table 4. This observation supports findings in the work [17] that demonstrated the benefits of predicting a probability distribution rather than a single point - several candidate locations shifted the single output in different directions. Similarly, the MOP-type architecture outperformed SOP-type in both Probabilistic and Geospatial variations, though the difference was relatively low.
@@ 4 Discussion and Conclusion
According to [12], the user-level geolocation is more effective in some cases, however, fine-grained analysis is possible only for the tweet-level predictions and thus the latter was chosen as the main subject of this study. Nonetheless, the problem of the user’s home location prediction was explored as one of the possible estimates based on the PMOP-type model predictions. Although this wasn’t the foremost performance measure, the results reveal a higher spatial accuracy on the per-user task than on the per-tweet measurements. Hence PMOP-type models are suggested for the user’s home location prediction as the comparable transformers-based solution.